OpenStudy (anonymous):

a body is thrown vertically upwards from a wall oh height 2.2m with a speed of 10m/s the distance travelled in the last second of its motion is (g=10m/s^2)

OpenStudy (ajprincess):

s=ut+1/2at^2 -2.2 =10t-1/2*10*t^2 50t^2-100t-22=0 25t^2-50t-11=0 t^2-2t-11/25=0 t^2-2t+1=11/25+1 (t-1)^2=36/25 t-1=+/-6/5 t=6/5+1 t=11/5 v=u+at 0=10-10t1 10=10t1 t1=1 t2=11/5-1 =6/5 d=0+10(6/5-1/2) =10((12-5)/10) =7